# The Analysis of Willmore Surfaces and its Generalizations in Higher Dimensions

### v1

@@ II.2.3 Analysis in high dimension
Given a closed orientable $n$-dimensional manifold $\Sigma$ and an immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{d}$ with $\mathrm{I\!I}_{\vec{\Phi}}\in L^{n}(\Sigma^{n})$, is it possible to construct counterexamples to the existence of controlled harmonic coordinates? Or can we construct controlled harmonic coordinates on a ”large” subset of $\Sigma$?
The next question is to know whether this atlas of harmonic coordinates can be ”uniformized” to obtain a global metric on $\Sigma$ conformal to $g_{\vec{\Phi}}$ together with some control following from an estimate of the form (II.7). A first attempt would be to find a metric having constant scalar curvature. This question is at the origin of the Yamabe problem and is not solvable in full generality, the topology of $\Sigma$ plays a role. An alternative has been proposed recently in [105], where the authors are able to solve locally $\Delta_{\bar{g}}\mathrm{Scal}_{\bar{g}}=0$ for some metric $\bar{g}$ conformal to $g$ in dimension 4. This equation is always solvable as long as the Sobolev inequalities are verified for the metric $g$ and the following regularity estimate holds
@@
Sobolev inequalities are a consequence of the isoperimetric inequality, which always hold for submanifolds having small $L^{n}$-norm of its mean curvature, thanks to Michael–Simon inequality [5, 135]. A global version of the equation $\Delta_{\bar{g}}\mathrm{Scal}_{\bar{g}}=0$ would be that the $Q$-curvature of $\bar{g}$ is pointwise controlled by $|\mathrm{Riem}^{\bar{g}}|^{2}_{\bar{g}}$.
@@
Given an immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{m}$ from a closed $4$-dimensional oriented manifold $\Sigma$ verifying (II.7), can we find a metric $\bar{g}$ conformal to $g_{\vec{\Phi}}$ such that the $Q$-curvature of $\bar{g}$ is constant? Is pointwise controlled by $|\vec{\mathrm{I\!I}}_{\vec{\Phi}}|^{4}_{g_{\vec{\Phi}}}$?
In order to know whether an equation on a manifold is regularizing or not, one needs to develop a refined PDE theory. In particular, one needs to control the constants involved in inequalities such as Sobolev or Poincaré inequalities and elliptic regularity. Such inequalities are directly related to isoperimetric problems, see for instance [33, Chapter IV]. Gallot [62] proved that the isoperimetric inequalities hold on a complete manifold $(\Sigma^{n},g)$ as soon as $\mathrm{Ric}_{g}\in L^{p}(\Sigma,g)$ for some $p>\frac{n}{2}$. Moreover Micheal–Simon inequality (see Brendle [26] for the optimal constants) implies that an isoperimetric inequality always holds for immersed $n$-dimensional submanifolds of arbitrary codimension in a Euclidean space whose mean curvature has small $L^{n}$-norm.
@@
Given a complete manifold $(\Sigma^{n},g)$, does there exists a number $s>0$ such that if $\mathrm{Riem}_{g}\in W^{s,\frac{n}{s+2}}(\Sigma,g)$, then the isoperimetric inequality holds?
@@
$\sup_{k\in\mathbb{N}}\left\|\vec{\mathrm{I\!I}}_{\vec{\Phi}_{k}}\right\|_{W^{s,\frac{n}{s+1}}(\Sigma,g_{\vec{\Phi}})}\leq C.$
@@
The Sobolev space $W^{s,\frac{n}{s+1}}(\Sigma^{n})$ always embeds in $L^{n}(\Sigma^{n})$. Hence, one should not expect a strong limit in full generality, but rather a weak limit on $\Sigma$ minus a set having small measure. The case $s=\frac{n-2}{2}$ and $n$ even has been treated in [129], based on the approximation result obtained in Theorem IV.23.
@@
Up to rescaling, we can assume that each $\vec{\Phi}_{k}(\Sigma)$ is contained in the unit ball $\mathbb{B}^{d}(0,1)$ and has diameter 1. Then, there exist a closed $C^{1}$ manifold $\Sigma_{\infty}$ (possibly non-connected) and a weak immersion $\vec{\Phi}_{\infty}\colon\Sigma_{\infty}\to\mathbb{R}^{d}$ such that the following holds. There exists a finite number of points $\vec{q}_{1},\ldots,\vec{q}_{K}\in\mathbb{R}^{d}$ such that for each $r>0$, we have the following convergence in the pointed Gromov–Hausdorff topology
@@ III.1 Notation
    For a real-valued function $f$, we denote $f^{+}\coloneqq\max(f,0)$, $f^{-}\coloneqq\max(-f,0)$.
@@
    We denote by $\mathcal{L}^{n}$ the Lebesgue measure on $\mathbb{R}^{n}$. For $U\subset\mathbb{R}^{n}$ with $\mathcal{L}^{n}(U)<\infty$, $f\in L^{1}(U)$, we write $\fint_{U}f\coloneqq(\mathcal{L}^{n}(U))^{-1}\int_{U}f$.
@@
    Let $(V,g)$ be an $n$-dimensional inner product space with a positively oriented orthonormal basis $(e_{1},\dots,e_{n})$. The Hodge star operator $_{g}$ is defined as the unique linear operator from $\bigwedge^{k}V$ to $\bigwedge^{n-k}V$ satisfying α∧*_g  β=⟨α,β⟩_g  e_1∧⋯∧e_n,   for all   α,β∈⋀^kV. See for instance [94, Section 3.3].
    When $V=\mathbb{R}^{n}$ and $g=g_{\text{std}}$ is the standard inner product on $\mathbb{R}^{n}$, we write $\star$ instead of $_{g}$.
@@
    Let $U\subset\mathbb{R}^{n}$ be an open set and let $g$ be a Riemannian metric on $U$. On the space of differential $\ell$-forms on $U$, we define the codifferential (see [94, Definition 3.3.1]) by
@@
    For $\vec{u}=(\vec{u}_{1},\vec{u}_{2})$, $\vec{v}=(\vec{v}_{1},\vec{v}_{2})$, where $\vec{u}_{i},\vec{v}_{i}\in\mathbb{R}^{3}$, denote →u⋅→v≔∑_i=1^2 →u_i ⋅→v_i,  →u×→v≔∑_i=1^2 →u_i ×→v_i.
@@
    Let $U\subset\mathbb{R}^{n}$ be open and $\vec{\Phi}\colon U\to\mathbb{R}^{m}$ be an immersion. Set $g=g_{\vec{\Phi}}=\vec{\Phi}^{*}g_{\text{std}}$, where $g_{\textup{std}}$ is the Euclidean metric on $\mathbb{R}^{m}$. Denote g_ij=∂_i→Φ⋅∂_j→Φ, (g^ij)=(g_ij)^-1, detg=det(g_ij), dvol_g=(detg)^12dx^1∧⋯∧dx^n. The pullback metric induces a pairing on $T^{*}U$ with $\langle dx^{i},dx^{j}\rangle_{g}=g^{ij}$; we denote by $|\cdot|_{g}$ the associated pointwise norm. We also write $|\cdot|_{\mathbb{R}^{m}}$ for the Euclidean norm in $\mathbb{R}^{m}$ (abbreviated to $|\cdot|$ when unambiguous).
@@
    $\vec{\mathrm{I\!I}}(X,Y)\coloneqq\pi_{\vec{n}}X(d\vec{\Phi}(Y)),\qquad\text{for all }X,Y\in T_{p}U\text{ and all }p\in U.$
@@ III.2 A useful formula
[168, Lemma 1.8] Let $\Sigma$ be a $2$-dimensional oriented smooth manifold, and $\vec{\Phi}\colon\Sigma\rightarrow\mathbb{R}^{3}$ be a smooth immersion. Let $_{g}$ be the Hodge star operator with respect to $g=g_{\vec{\Phi}}$. Then we have
@@
For any $k\in\{1,2\}$, note
$\partial_{k}(\vec{n}\cdot\vec{n})=2\,\partial_{k}\vec{n}\cdot\vec{n}=0,$
hence
@@
$(\vec{u}\times\vec{v})\times\vec{w}=(\vec{u}\cdot\vec{w})\vec{v}-(\vec{v}\cdot\vec{w})\vec{u}$
for all $\vec{u},\vec{v},\vec{w}\in\mathbb{R}^{3}$. By (III.11), we also get
$dx_{s}\wedge d\vec{n}$ $=\varepsilon^{sk}\,\partial_{k}n\,\alpha=-g^{ij}\epsilon^{sk}\,\mathrm{I\!I}_{kj}\,\partial_{i}\vec{\Phi}\,\alpha,$
Therefore,
@@
where the second equality follows from the identity
$f(s)\varepsilon^{ki}+f(k)\varepsilon^{is}+f(i)\varepsilon^{sk}=0$
which holds for any $i,k,s\in\{1,2\}$ and real-valued function $f$ defined on $\{1,2\}$. Since (III.12) holds for any $s\in\{1,2\}$, (III.10) follows immediately. ∎
@@ III.3 Some estimates for elliptic equations
Lemma III.2 ([77, Lemma 4.12]).
Let $n\in\mathbb{N}^{+}$. Suppose $\{a^{ij}\}_{i,j=1}^{n}\subset L^{\infty}(B^{n})$ satisfies
$\Lambda^{-1}|\xi|^{2}\leq a^{ij}(x)\,\xi_{i}\,\xi_{j}\leq\Lambda|\xi|^{2},\qquad\text{for a.e. }x\in B^{n}\text{ and all }\,\xi\in\mathbb{R}^{n},$ (III.13)
where $\Lambda>0$ is a constant. Suppose $u\in W^{1,2}(B^{n})$ is a weak solution of the equation $\partial_{j}(a^{ij}\,\partial_{i}u)=0$ in $B^{n}$, i.e.
$\int_{B^{n}}a^{ij}\,\partial_{i}u\,\partial_{j}\varphi=0,\qquad\text{for all }\,\varphi\in C^{\infty}_{c}(B^{n}).$
@@
$\begin{cases}\partial_{i}(a^{ij}\,\partial_{j}\phi)=\partial_{1}u\,\partial_{2}v-\partial_{2}u\,\partial_{1}v&\mbox{in }D^{2},\\[5.69054pt] \phi=0&\mbox{on }\partial D^{2},\end{cases}$
and we also have $\phi\in C(\overline{D^{2}})$ with the estimate
@@
We have a stronger version of this result, see Proposition IV.8.
@@ III.4 Bourgain–Brezis inequality
where $C=C(n)$, $\mathbb{T}^{n}$ is the $n$-dimensional torus. In the case $n=2$, F. Da Lio and T. Rivière [48] provided a proof for a stronger version of (III.14) in 2020. Here we denote $\dot{W}^{-1,2}(\mathbb{R}^{2})$ as the dual space of the homogeneous Sobolev space $\dot{W}^{1,2}(\mathbb{R}^{2})$.
@@
Assume $g\in L^{1}(\mathbb{R}^{2},\mathbb{R}^{2})$, $f\in\dot{W}^{-1,2}(\mathbb{R}^{2},\mathbb{R}^{2})$, $u$ be a tempered distribution on $\mathbb{R}^{2}$ such that $\nabla u=f+g$. Then there exists $c\in\mathbb{R}$, and a universal constant $C$ such that
@@ III.5 Quasiconformal maps and Beltrami equations
Lemma III.7 ([90, Theorem 4.30]).
Let $\mu\in L^{\infty}(\mathbb{C},\mathbb{C})$ satisfies $\|\mu\|_{L^{\infty}(\mathbb{C})}<1$. Then there exists a homeomorphism $f\colon\bar{\mathbb{C}}\rightarrow\bar{\mathbb{C}}$ which is a quasiconformal map of $\mathbb{C}$ and satisfies the Beltrami equation $\partial_{\bar{z}}f=\mu\,\partial_{z}f$. Such $f$ is uniquely determined by the conditions $f(0)=0$, $f(1)=1$, $f(\infty)=f(\infty)$.
@@ III.6 Hardy space $\mathcal{H}^{1}$
Standard singular integral theory (see for instance [182, Theorem III.4]) implies the following results.
@@ III.7 Functions of VMO
Then by induction we have
$\left|\fint_{B}f-\fint_{2\tilde{B}}f\right|\leq C(n,K)\,i\beta,\qquad\text{for all }\,i\geq 0\,\text{ and all }B\in\mathcal{G}_{i}.$
@@
Lemma III.14 ([162, Theorem 2], [92, Section 3]).
@@
$\beta_{r,\varphi(U)}(f\circ\varphi^{-1})\leq C(K)\,\beta_{r,U}(f)\qquad\text{for all }\,r>0.$
@@
$\beta_{r,\mathbb{R}^{n}}(\tilde{f})\leq C(U)\big(\beta_{r,U}(f)+r\|f\|_{L^{\infty}(U)}\big),\quad\text{for all }\,r>0.$

### v3

@@ II.2.3 Analysis in high dimensions
Given a closed orientable $n$-dimensional manifold $\Sigma$ and an immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{d}$ with $\mathrm{I\!I}_{\vec{\Phi}}\in L^{n}(\Sigma^{n})$, is it possible to construct counterexamples to the existence of controlled harmonic coordinates? Or can we construct controlled harmonic coordinates on a “large” subset of $\Sigma$?
The next question is to know whether this atlas of harmonic coordinates can be “uniformized” to obtain a global metric on $\Sigma$ conformal to $g_{\vec{\Phi}}$ together with some control following from an estimate of the form (II.7). A first attempt would be to find a metric having constant scalar curvature. This question is at the origin of the Yamabe problem and is not solvable in full generality, the topology of $\Sigma$ plays a role when it has a boundary. An alternative has been proposed recently in [108], where the authors are able to solve locally $\Delta_{\bar{g}}\mathrm{Scal}_{\bar{g}}=0$ for some metric $\bar{g}$ conformal to $g$ in dimension 4. This equation is always solvable as long as the Sobolev inequalities are verified for the metric $g$ and the following regularity estimate holds
@@
Sobolev inequalities are a consequence of the isoperimetric inequality, which always holds for submanifolds with a small $L^{n}$-norm of its mean curvature, thanks to the Michael–Simon inequality [5, 139]. A global version of the equation $\Delta_{\bar{g}}\mathrm{Scal}_{\bar{g}}=0$ would be that the $Q$-curvature of $\bar{g}$ is pointwise controlled by $|\mathrm{Riem}^{\bar{g}}|^{2}_{\bar{g}}$.
@@
Given an immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{m}$ from a closed $4$-dimensional oriented manifold $\Sigma$ verifying (II.7), can we find a metric $\bar{g}$ conformal to $g_{\vec{\Phi}}$ such that the $Q$-curvature of $\bar{g}$ is constant? Is it pointwise controlled by $|\vec{\mathrm{I\!I}}_{\vec{\Phi}}|^{4}_{g_{\vec{\Phi}}}$?
In order to know whether an equation on a manifold is regularizing or not, one needs to develop a refined PDE theory. In particular, one needs to control the constants involved in inequalities such as Sobolev or Poincaré inequalities and elliptic regularity. Such inequalities are directly related to isoperimetric problems, see for instance [33, Chapter IV]. Gallot [62] proved that the isoperimetric inequality holds on a complete manifold $(\Sigma^{n},g)$ as soon as $\mathrm{Ric}_{g}\in L^{p}(\Sigma,g)$ for some $p>\frac{n}{2}$. Moreover, the Michael–Simon inequality (see Brendle [26] for the optimal constants) implies that an isoperimetric inequality always holds for immersed $n$-dimensional submanifolds of arbitrary codimension in a Euclidean space whose mean curvature has small $L^{n}$-norm.
@@
Given a complete manifold $(\Sigma^{n},g)$, is there a number $s>0$ such that if $\mathrm{Riem}_{g}\in W^{s,\frac{n}{s+2}}(\Sigma,g)$, then the isoperimetric inequality holds?
@@
$\sup_{k\in\mathbb{N}}\left\|\vec{\mathrm{I\!I}}_{\vec{\Phi}_{k}}\right\|_{W^{s,\frac{n}{s+1}}(\Sigma,g_{\vec{\Phi}})}<+\infty.$
@@
The Sobolev space $W^{s,\frac{n}{s+1}}(\Sigma^{n})$ always embeds in $L^{n}(\Sigma^{n})$. Hence, one should not expect a strong limit in full generality, but rather a weak limit on $\Sigma$ minus a set having small measure. The case $s=\frac{n-2}{2}$ and $n\geq 4$ even has been treated in [133], based on the approximation result obtained in Theorem IV.23.
@@
Up to rescaling, we can assume that each $\vec{\Phi}_{k}(\Sigma)$ is contained in the unit ball $\mathbb{B}^{d}(0,1)$ and has diameter 1. Then, there exists a closed $C^{1}$ manifold $\Sigma_{\infty}$ (possibly non-connected) and a weak immersion $\vec{\Phi}_{\infty}\colon\Sigma_{\infty}\to\mathbb{R}^{d}$ such that the following holds. There exists a finite number of points $\vec{q}_{1},\ldots,\vec{q}_{K}\in\mathbb{R}^{d}$ such that for each $r>0$, we have the following convergence in the pointed Gromov–Hausdorff topology
@@ III.1 Notation
    For a real-valued function $f$, we denote $f^{+}\coloneqq\max(f,0)$ and $f^{-}\coloneqq\max(-f,0)$.
@@
    We denote by $\mathcal{L}^{n}$ the Lebesgue measure on $\mathbb{R}^{n}$. For $U\subset\mathbb{R}^{n}$ with $\mathcal{L}^{n}(U)<\infty$ and $f\in L^{1}(U)$, we write $\fint_{U}f\coloneqq(\mathcal{L}^{n}(U))^{-1}\int_{U}f$.
@@
    Let $(V,g)$ be an $n$-dimensional inner product space with a positively oriented orthonormal basis $(e_{1},\dots,e_{n})$. The Hodge star operator $_{g}$ is defined as the unique linear operator from $\bigwedge^{k}V$ to $\bigwedge^{n-k}V$ satisfying α∧*_g  β=⟨α,β⟩_g  e_1∧⋯∧e_n,   for all α,β∈⋀^kV. See for instance [95, Section 3.3]. When $V=\mathbb{R}^{n}$ and $g=g_{\text{std}}$ is the standard inner product on $\mathbb{R}^{n}$, we write $\star$ instead of $_{g}$.
@@
    Let $U\subset\mathbb{R}^{n}$ be an open set and let $g$ be a Riemannian metric on $U$. On the space of differential $\ell$-forms on $U$, we define the codifferential (see [95, Definition 3.3.1]) by
@@
    For $\vec{u}=(\vec{u}_{1},\vec{u}_{2})$, $\vec{v}=(\vec{v}_{1},\vec{v}_{2})$, where $\vec{u}_{i},\vec{v}_{i}\in\mathbb{R}^{3}$, we denote →u⋅→v≔∑_i=1^2 →u_i ⋅→v_i,  →u×→v≔∑_i=1^2 →u_i ×→v_i.
@@
    Let $U\subset\mathbb{R}^{n}$ be open and $\vec{\Phi}\colon U\to\mathbb{R}^{m}$ be an immersion. Set $g=g_{\vec{\Phi}}=\vec{\Phi}^{*}g_{\text{std}}$, where $g_{\textup{std}}$ is the Euclidean metric on $\mathbb{R}^{m}$. We denote g_ij=∂_i→Φ⋅∂_j→Φ, (g^ij)=(g_ij)^-1, detg=det(g_ij), dvol_g=(detg)^12dx^1∧⋯∧dx^n. The pullback metric induces a pairing on $T^{*}U$ with $\langle dx^{i},dx^{j}\rangle_{g}=g^{ij}$. We denote by $|\cdot|_{g}$ the associated pointwise norm. We also write $|\cdot|_{\mathbb{R}^{m}}$ for the Euclidean norm in $\mathbb{R}^{m}$ (abbreviated to $|\cdot|$ when unambiguous).
@@
    $\vec{\mathrm{I\!I}}(X,Y)\coloneqq\pi_{\vec{n}}X(d\vec{\Phi}(Y)),\qquad\forall X,Y\in T_{p}U,\ \forall p\in U.$
@@ III.2 A useful formula
[171, Lemma 1.8] Let $\Sigma$ be a $2$-dimensional oriented smooth manifold, and $\vec{\Phi}\colon\Sigma\rightarrow\mathbb{R}^{3}$ be a smooth immersion. Let $_{g}$ be the Hodge star operator with respect to $g=g_{\vec{\Phi}}$. Then we have
@@
For any $k\in\{1,2\}$, we have
$\partial_{k}(\vec{n}\cdot\vec{n})=2\,\partial_{k}\vec{n}\cdot\vec{n}=0.$
Hence, it holds
@@
$\forall,\vec{u},\vec{v},\vec{w}\in\mathbb{R}^{3},\qquad(\vec{u}\times\vec{v})\times\vec{w}=(\vec{u}\cdot\vec{w})\vec{v}-(\vec{v}\cdot\vec{w})\vec{u}.$
By (III.11), we also get
$dx_{s}\wedge d\vec{n}$ $=\varepsilon^{sk}\,\partial_{k}n\,\alpha=-g^{ij}\epsilon^{sk}\,\mathrm{I\!I}_{kj}\,\partial_{i}\vec{\Phi}\,\alpha.$
Therefore, it holds
@@
where the second equality follows from the identity, valid for any $i,k,s\in\{1,2\}$ and real-valued function $f$ defined on $\{1,2\}$
$f(s)\varepsilon^{ki}+f(k)\varepsilon^{is}+f(i)\varepsilon^{sk}=0.$
Since (III.12) holds for any $s\in\{1,2\}$, the relation (III.10) follows immediately. ∎
@@ III.3 Some estimates for elliptic equations
Lemma III.2 ([78, Lemma 4.12]).
Let $n\in\mathbb{N}^{+}$ and $\Lambda>0$. Suppose $\{a^{ij}\}_{i,j=1}^{n}\subset L^{\infty}(B^{n})$ satisfies
$\Lambda^{-1}|\xi|^{2}\leq a^{ij}(x)\,\xi_{i}\,\xi_{j}\leq\Lambda|\xi|^{2},\qquad\text{for a.e. }x\in B^{n}\text{ and all }\,\xi\in\mathbb{R}^{n}.$ (III.13)
Suppose $u\in W^{1,2}(B^{n})$ is a weak solution of the equation $\partial_{j}(a^{ij}\,\partial_{i}u)=0$ in $B^{n}$, i.e.
$\int_{B^{n}}a^{ij}\,\partial_{i}u\,\partial_{j}\varphi=0,\qquad\forall\varphi\in C^{\infty}_{c}(B^{n}).$
@@
$\begin{cases}\partial_{i}(a^{ij}\,\partial_{j}\phi)=\partial_{1}u\,\partial_{2}v-\partial_{2}u\,\partial_{1}v&\mbox{in }D^{2},\\[5.69054pt] \phi=0&\mbox{on }\partial D^{2}.\end{cases}$
We have $\phi\in C(\overline{D^{2}})$ with the estimate
@@
We prove a stronger version of this result in Proposition IV.8 below.
@@ III.4 Bourgain–Brezis inequality
where $C=C(n)$, $\mathbb{T}^{n}$ is the $n$-dimensional torus. In the case $n=2$, F. Da Lio and the third author [48] provided a proof for a stronger version of (III.14) in 2020. Here we denote $\dot{W}^{-1,2}(\mathbb{R}^{2})$ as the dual space of the homogeneous Sobolev space $\dot{W}^{1,2}(\mathbb{R}^{2})$.
@@
Assume $g\in L^{1}(\mathbb{R}^{2},\mathbb{R}^{2})$, $f\in\dot{W}^{-1,2}(\mathbb{R}^{2},\mathbb{R}^{2})$, $u$ be a tempered distribution on $\mathbb{R}^{2}$ such that $\nabla u=f+g$. Then there exist $c\in\mathbb{R}$ and a universal constant $C$ such that
@@ III.5 Quasiconformal maps and Beltrami equations
Lemma III.7 ([91, Theorem 4.30]).
Let $\mu\in L^{\infty}(\mathbb{C},\mathbb{C})$ be such that $\|\mu\|_{L^{\infty}(\mathbb{C})}<1$. Then there exists a quasiconformal homeomorphism $f\colon\bar{\mathbb{C}}\rightarrow\bar{\mathbb{C}}$ verifying the Beltrami equation $\partial_{\bar{z}}f=\mu\,\partial_{z}f$. Such $f$ is uniquely determined by the conditions $f(0)=0$, $f(1)=1$, $f(\infty)=f(\infty)$.
@@ III.6 Hardy space $\mathcal{H}^{1}$
Standard singular integral theory (see for instance [185, Theorem III.4]) implies the following results.
@@ III.7 Functions of VMO
Then by induction we have for all $i\geq 0$ and all $B\in\mathcal{G}_{i}$,
$\left|\fint_{B}f-\fint_{2\tilde{B}}f\right|\leq C(n,K)\,i\beta.$
@@
Lemma III.14 ([165, Theorem 2], [93, Section 3]).
@@
$\forall r>0,\qquad\beta_{r,\varphi(U)}(f\circ\varphi^{-1})\leq C(K)\,\beta_{r,U}(f).$
@@
$\forall r>0,\qquad\beta_{r,\mathbb{R}^{n}}(\tilde{f})\leq C(U)\big(\beta_{r,U}(f)+r\|f\|_{L^{\infty}(U)}\big).$